OpenStudy (anonymous):

A line passes through (3, –2) and (6, 2). Write an equation for the line in point-slope form. Rewrite the equation in standard form using integers. y – 3 = 4/3(x + 2); –4x + 3y = 17 y + 2 =4/3(x – 3); –4x + 3y = –18 y + 2 =4/3 (x + 3); –4x + 3y = –18 y – 2 =4/3(x – 3); –4x + 3y = 18

OpenStudy (anonymous):

slope is 2+2/6-3 = 4/3

OpenStudy (anonymous):

so you have y + 2 = 4/3(x-3) or y-2 = 4/3(x-6)

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

its the first one, both are correct, but you're given the first one in the options